The speed of laserlight in a transparent solid is23 and 24 could you check my answers
Answer #1

Q.23) Speed of light in the solid = 0.49c and Speed of light in vacuum = c

where c = 3 x 108 m/s

So the refractive index of the solid will be: n = c/v = c/0.49c = 2.0408

Now for Critical Angle, the incident medium is the solid and the refractive medium is the vacuum.

And the relation between the indices of refraction and the critical angle is given by:

\small sin\theta_c = \frac{n_r}{n_i}

=> \small \theta_c = sin^{-1}(\frac{n_r}{n_i}) = sin^{-1}(\frac{1}{2.0408}) = 29.34\degree \approx 29\degree

Q.24) True. Our pupils (black hole in the eye) get dilated when in dark to absorb as much light (hence as much information of the surrounding) as we can. If the surrounding is well lit then our pupils contract and make a fine pinhole. This reduces the aberrations since only the paraxial rays (light rays making small angle with the optic axis) are reaching your eyes which will effectively describe the image accurately.